OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 4.5 hours with a 45 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. None of these systems can be used to solve this problem. d = 7(r + 45) d = 4.5(r + 45) and d = 2.5(r − 45) d = 4.5(r − 45) and d = 2.5(r + 45)

OpenStudy (anonymous):

d = 4.5(r − 45) and d = 2.5(r + 45) or 4.5 (r - 45) = 2.5 (r + 45)